One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by checkley75(3666)

X+12=Y, X=Y-12
X+2Y=39 NOW SUBSTITUTE (Y-12) FOR X IN THIS EQUATION & SOLVE FOR X
(Y-12)+2Y+39
Y-12+2Y=39
3Y=39+12
3Y=51
Y=51/3
Y=17 ANSWER FOR THE LARGER NUMBER.
X+12=17
X=17-12
X=5 ANSWER FOR THE SMALLER NUMBER.
PROOF
5+2*17=39
5+34
39=39